OpenStudy (anonymous):

Find f(-3) of the function, f(x) = {x^3 + 2x^2 - x + 6, x < -3 |-3x - 3|, x >= -3

zepdrix (zepdrix):

Hey frog :) \[\Large f(x)=\cases{x^3+2x^2-x+6, &$x\lt-3$\\ |-3x-3|, &$x\ge-3$}\]

zepdrix (zepdrix):

With a piece-wise function we only use ONE PIECE at a time. So based on which x value they give us, that will tell us which part to use. Example: If we wanted to evaluate this function at -4,\[\Large f(x)=\cases{\color{orangered}{x^3+2x^2-x+6,} &$x\lt-3$\\ |-3x-3|, &$x\ge-3$}\]We would use this piece, `because this is the piece we use when our x value is less than -3`

zepdrix (zepdrix):

So in that example we would get:\[\Large\rm f(-4)=(-4)^3+2(-4)^2-(-4)+6\]Which simplifies to something..

zepdrix (zepdrix):

If we want to evaluate this function at -3, which piece do you think we should use? Which part has an "or equal to" for the point -3?

OpenStudy (anonymous):

The second part?

OpenStudy (anonymous):

So: |-3(-3) - 3| = |9 - 3| = |6| = 6 ?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Yayyyy good job \c:/

OpenStudy (anonymous):

thanks!